# access point ADAS advanced driver-assistance systems

### v1

@@ Appendix B Proof of Lemma 2
The conditional success probability is derived as
$P_{\rm d}=\mathbb{P}\left(\xi(R)\geq\gamma\;|\;\Phi\right)$
$=\mathbb{P}\left[\frac{\xi_{0}{R}^{-2\alpha}\sigma_{\rm c}}{1+\xi_{0}\sum_{{\bf w}\in\Phi^{\prime}}{h}_{\bf w}d_{\bf w}^{-\alpha}\textbf{1}({\bf w}\in\mathcal{C})}\geq\gamma\;|\;\Phi\right]$
=$\overset{(a)}{=}\mathbb{E}_{h_{\bf w}}\left[\exp\left(\frac{-\gamma-\gamma\xi_{0}\sum_{{\bf w}\in\Phi^{\prime}}{h}_{\bf w}d_{\bf w}^{-\alpha}\textbf{1}({\bf w}\in\mathcal{C})}{\xi_{0}R^{-2\alpha}}\right)\right]$
$=e^{\left(\frac{-\gamma}{\xi_{0}R^{-2\alpha}}\right)}\mathbb{E}_{h_{\bf w}}\left[\exp\left(\frac{-\gamma\xi_{0}\sum_{{\bf w}\in\Phi^{\prime}}{h}_{\bf w}d_{\bf w}^{-\alpha}\textbf{1}({\bf w}\in\mathcal{C})}{\xi_{0}R^{-2\alpha}}\right)\right]$
$=e^{\left(\frac{-\gamma}{\xi_{0}R^{-2\alpha}}\right)}\left(\prod_{{\bf w}\in\Phi^{\prime}}\mathbb{E}_{h_{\bf w}}\exp\left(\frac{-\gamma\xi_{0}d_{\bf w}^{-\alpha}{h}_{w}}{\xi_{0}R^{-2\alpha}}\right)\right)$
=$\overset{(b)}{=}\exp{\left(\frac{-\gamma}{\xi_{0}R^{-2\alpha}}\right)}\left(\prod_{{\bf w}\in\Phi^{\prime}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right).$
Here $\mathcal{C}$ denotes the active radar set in $\phi$ (a realization of $\Phi$). Step (a) is due to the exponential distribution of $\sigma_{\rm c}$. Step (b) follows from the Laplace transform of the exponentially distributed $h_{\bf w}$.
@@ Appendix C Proof of Lemma 3
$\mathbb{E}\left[P^{l}_{\rm d}\right]$ $=\mathbb{E}\left[\exp{\left(\frac{-\gamma}{\xi_{0}R^{-2\alpha}}\right)}\left(\prod_{{\bf w}\in\Phi^{\prime}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)\right]$
$=\exp{\left(\frac{-l\gamma}{\xi_{0}R^{-2\alpha}}\right)}\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)\right].$ (7)
@@
$\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)\right]=$ $\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}_{0}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)\right]\cdot$
$\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}\backslash\Phi^{\prime}_{0}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)\right],$ (8)
@@
$\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}_{0}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)\right]=\exp\left(-\lambda\delta\int_{0}^{\infty}1-\mathcal{T}_{z}^{l}{\rm d}z\right),$ (9)
@@
$\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}\backslash\Phi^{\prime}_{0}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)\right]=$
$\mathbb{E}_{\Phi_{I}}\left[\mathbb{E}_{\Phi^{\prime}_{i}}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}_{i}\backslash\Phi^{\prime}_{0}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)\right]\bigg{|}\Phi_{I}\right]$
$=\mathbb{E}_{\Phi_{I}}\left[\exp\left(-\delta\lambda\int_{a_{i}}^{b_{i}}1-\frac{R^{-2\alpha}}{R^{-2\alpha}+\gamma w(x,z,\theta))^{-\alpha}}{\rm d}z\right)\right]$

### v2

@@ Appendix B Proof of Lemma 2
The conditional success probability is defined as the conditional CDF of $\xi(R)$, i.e.,
$P_{\rm d}=\mathbb{P}\left(\xi(R)\geq\gamma\;|\;\Phi\right)$ (12)
$=\mathbb{P}\left[\frac{\xi_{0}{R}^{-2\alpha}\sigma_{\rm c}}{1+\xi_{0}\sum_{{\bf w}\in\Phi^{\prime}}{h}_{\bf w}d_{\bf w}^{-\alpha}\textbf{1}({\bf w}\in\mathcal{C})}\geq\gamma\;|\;\Phi\right].$ (13)
Here $\mathcal{C}$ denotes the active radar set in $\phi$ (a realization of $\Phi$). Now, due to the Swerling-I model for the fluctuating cross-section, $\sigma_{\rm c}$ is exponentially distributed with mean $\bar{\sigma}$. Accordingly, we calculate the above as
$P_{\rm d}$ $=\mathbb{P}\left[\sigma_{\rm c}\geq\frac{\gamma\left(1+\xi_{0}\sum_{{\bf w}\in\Phi^{\prime}}{h}_{\bf w}d_{\bf w}^{-\alpha}\textbf{1}({\bf w}\in\mathcal{C})\right)}{\xi_{0}R^{-2\alpha}}\;|\;\Phi\right]$
$\overset{(a)}{=}\mathbb{E}_{h_{\bf w}}\left[\exp\left(\frac{-\gamma-\gamma\xi_{0}\sum_{{\bf w}\in\Phi^{\prime}}{h}_{\bf w}d_{\bf w}^{-\alpha}\textbf{1}({\bf w}\in\mathcal{C})}{\bar{\sigma}\xi_{0}R^{-2\alpha}}\right)\right].$
In the above, the step (a) is due to the fact that $\sigma_{\rm c}$ is exponentially distributed. Due to the conditioning on $\Phi$, the locations of the interferers, i.e., $\bf w$} is fixed and not random. The only randomness that remains in the resulting expression is due to the fast-fading $h_{\bf w}$, with respect to which we take the expectation in the remaining steps. This is evaluated as
$P_{\rm d}$ $=e^{\left(\frac{-\gamma}{\bar{\sigma}\xi_{0}R^{-2\alpha}}\right)}\mathbb{E}_{h_{\bf w}}\left[\exp\left(\frac{-\gamma\xi_{0}\sum_{{\bf w}\in\Phi^{\prime}}{h}_{\bf w}d_{\bf w}^{-\alpha}\textbf{1}({\bf w}\in\mathcal{C})}{\bar{\sigma}\xi_{0}R^{-2\alpha}}\right)\right]$
$=e^{\left(\frac{-\gamma}{\bar{\sigma}\xi_{0}R^{-2\alpha}}\right)}\left(\prod_{{\bf w}\in\Phi^{\prime}}\mathbb{E}_{h_{\bf w}}\exp\left(\frac{-\gamma\xi_{0}d_{\bf w}^{-\alpha}{h}_{w}}{\bar{\sigma}\xi_{0}R^{-2\alpha}}\right)\right)$
$\overset{(b)}{=}\underbrace{\exp{\left(\frac{-\gamma}{\xi_{0}R^{-2\alpha}}\right)}}_{\text{Impact of noise}}\underbrace{\left(\prod_{{\bf w}\in\phi^{\prime}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)}_{\text{Impact of interference}}.$
Step (b) follows from the Laplace transform of the exponentially distributed $h_{\bf w}$.
@@ Appendix C Proof of Lemma 3
$\mathbb{E}\left[P^{l}_{\rm d}\right]$ $=\mathbb{E}\left[\exp{\left(\frac{-l\gamma}{\xi_{0}R^{-2\alpha}}\right)}\left(\prod_{{\bf w}\in\Phi^{\prime}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)^{l}\right]$
$=\exp{\left(\frac{-l\gamma}{\xi_{0}R^{-2\alpha}}\right)}\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)^{l}\right].$ (14)
@@
$\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)^{l}\right]=$ $\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}_{0}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)^{l}\right]\cdot$
$\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}\backslash\Phi^{\prime}_{0}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)^{l}\right],$ (15)
@@
$\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}_{0}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)^{l}\right]=\exp\left(-\lambda\delta\int_{0}^{\infty}1-\mathcal{T}_{z}^{l}{\rm d}z\right),$ (16)
@@
$\mathbb{E}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}\backslash\Phi^{\prime}_{0}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)^{l}\right]=$
$\mathbb{E}_{\Phi_{I}}\left[\mathbb{E}_{\Phi^{\prime}_{i}}\left[\left(\prod_{{\bf w}\in\Phi^{\prime}_{i}\backslash\Phi^{\prime}_{0}}\frac{1}{1+\frac{\gamma d_{\bf w}^{-\alpha}}{R^{-2\alpha}}}\right)^{l}\right]\bigg|\Phi_{I}\right]$
$=\mathbb{E}_{\Phi_{I}}\left[\exp\left(-\delta\lambda\int_{a_{i}}^{b_{i}}1-\left(\frac{R^{-2\alpha}}{R^{-2\alpha}+\gamma w(x,z,\theta))^{-\alpha}}\right)^{l}{\rm d}z\right)\right]$
@@ Appendix D Proof of Theorem 7
Given that in each slot the detection success probability is $P_{\rm d}$, we are interested in the distribution of the run length of detection successes. This problem is solved using generating functions. In particular, the combinatorial class of all sequences without a run of $m$ successes (S), i.e., $\xi\geq\gamma$ in a row can be written as
$\sum_{k\geq 0}(\mbox{length}_{<m}(S)\,F)^{k}\,\mbox{seq}_{<m}(S),$
where $F$ denotes the failure, i.e., $\xi<\gamma$, with corresponding counting generating function
$L(s,f)={\sum_{0\leq j<m}s^{j}\over 1-(\sum_{0\leq j<m}s^{j})t}={1-s^{m}\over 1-s-(1-s^{m})f}.$
We introduce the conditional success probability by replacing $s$ with $P_{\rm d}$ and $f$ with $q$, where $q=1-P_{\rm d}$. The generating functional in terms of $s$ thus takes the form
$G(s)={1-P_{\rm d}^{m}s^{m}\over 1-s+P_{\rm d}^{m}s^{m+1}(1-P_{\rm d}))}.$
The coefficient of $s^{n}$ in $G(s)$ is $\mathbb{P}(L<m).$ Note that the function $1/(1-s(1-P_{\rm d}^{m}s^{m}(1-P_{\rm d})))$ can be rewritten as
$\sum_{k\geq 0}s^{k}(1-P_{\rm d}^{m}s^{m}(1-P_{\rm d}))^{k}$
$=\sum_{k\geq 0}\sum_{j\geq 0}{k\choose j}(-P_{\rm d}^{m}(1-P_{\rm d}))^{j}s^{k+jm}.$ (17)
The coefficient of $s^{n}$ in this function is $c(n)=\sum_{j\geq 0}{n-jm\choose j}(-P_{\rm d}^{m}(1-P_{\rm d}))^{j}$. Therefore the coefficient of $s^{n}$ in $G(s)$ is $c(n)-P_{\rm d}^{m}c(n-m)$. Finally,
$\mathbb{P}(L\geq m)=1-\mathbb{P}(L<m)$
$=$ $P_{\rm d}^{m}c(n-m)+1-c(n)$
$=$ $P_{\rm d}^{m}\sum_{j\geq 0}(-1)^{j}{n-(j+1)m\choose j}(P_{\rm d}^{m}(1-P_{\rm d}))^{j}+$
$\sum_{j\geq 1}(-1)^{j+1}{n-jm\choose j}(P_{\rm d}^{m}(1-P_{\rm d}))^{j}$
$=$ $\sum_{j\geq 1}(-1)^{j+1}\left[{n-jm\choose j-1}+{n-jm\choose j}(1-P_{\rm d})\right]$
$P_{\rm d}^{jm}(1-P_{\rm d})^{j-1}$
$=$ $\sum_{j\geq 1}(-1)^{j+1}\left[{n-jm\choose j-1}P_{\rm d}+{n-jm+1\choose j}(1-P_{\rm d})\right]$
$P_{\rm d}^{jm}(1-P_{\rm d})^{j-1}$
$=$ $\sum_{j\geq 1}(-1)^{j+1}\left[P_{\rm d}+{n-jm+1\over j}\,(1-P_{\rm d})\right]$
${n-jm\choose j-1}\,P_{\rm d}^{jm}(1-P_{\rm d})^{j-1}.$
Appendix E Proof of Theorem 2